[figure: 9.7 cm]

EFFECT OF HEAVY ION IRRADIATION ON THE MAGNETIC

PROPERTIES OF CuGeO_{3}

L. Gladczuk^{a}, E. Mosiniewicz-Szablewska^{a}, R. Szymczak^{a}, M. Baran^{a}, H.A. Dabkowska^{b}, J. Choinski^{c}, T. Czosnyka^{c} and H. Szymczak^{a} aInstitute of Physics, Polish Academy of Sciences, Warsaw, Poland

bBrockhouse Institute for Materials Research, McMaster University, Hamilton, Canada cHeavy Ion Laboratory, Warsaw University, Warsaw, Poland

The influence of heavy ion irradiation on the magnetic properties of the spin-Peierls compound CuGeO_{3} has been investigated by means of the electron paramagnetic resonance (EPR), the spontaneous magnetization and the low field ac susceptibility. The measurements were performed on the CuGeO_{3} single crystals before and after irradiation with 80 MeV oxygen ions in the wide temperature range 2-300 K. It was found that the irradiation induced an increase in the EPR signal intensity, magnetization and susceptibility of this material. It was also shown that the irradiation do not lead to a decrease of the spin-Peierls transition temperature and to the formation of an antiferromagnetic phase which is expected to appear in disordered spin-Peierls systems. The analysis of the experimental data in the spin-Peierls phase indicates a complete destruction of both the energy gap and the dimerization within the Cu chains.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Leszek Gladczuk

Address for correspondence : Instytut Fizyki PAN Al. Lotników 32/46 02-668 Warszawa

Email address : gladl@ifpan.edu.pl